OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95. (2 points) Question 1 options: 1) 203 2) 197 3) 168 4) 163

jimthompson5910 (jim_thompson5910):

Are you familiar with this formula? \[\Large a_{n} = a_1 + d(n-1)\]

OpenStudy (anonymous):

im confused @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Plug in n = 18. Then plug in the given values for a1 and a18. Finally solve for d \[\Large a_{n} = a_1 + d(n-1)\] \[\Large a_{18} = a_1 + d(18-1)\] \[\Large 95 = -7 + d(18-1)\] solve for d to get d = ???

OpenStudy (anonymous):

what is the value for D im confused

jimthompson5910 (jim_thompson5910):

you have to isolate d

jimthompson5910 (jim_thompson5910):

\[\Large 95 = -7 + d(18-1)\] \[\Large 95 = -7 + d(17)\] \[\Large 95 = -7 + 17d\] \[\Large d = ???\]

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

95

jimthompson5910 (jim_thompson5910):

so the nth term is... \[\Large a_{n} = a_1 + d(n-1)\] \[\Large a_{n} = -7 + d(n-1)\] \[\Large a_{n} = -7 + 6(n-1)\] \[\Large a_{n} = -7 + 6n-6\] \[\Large a_{n} = 6n-13\]

OpenStudy (anonymous):

95

jimthompson5910 (jim_thompson5910):

Plug in n = 35 to compute the 35th term \[\Large a_{n} = 6n-13\] \[\Large a_{35} = 6(35)-13\] \[\Large a_{35} = ??\]

OpenStudy (anonymous):

197

jimthompson5910 (jim_thompson5910):

it's not 95. Try again

jimthompson5910 (jim_thompson5910):

yeah it's 197

OpenStudy (anonymous):

thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

can you help me with another problem? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok